Accuracy in forecasting can be measured by:
A) MSE B) MRP C) MAPE D) MTM E) A & C
Answer #1

The correct answer is option \(\mathrm{E}\).

Both option A. and option C. are commonly used measures for summarizing historical errors.

MSE - Mean Squared Error (Average of squared error) MAPE - Mean Absolute Percentage Error (Average of percentage error)

$$ \begin{aligned} &\operatorname{MSE}=\frac{\Sigma\left(\text { Actual }_{t}-\text { Forecast }_{t}\right)^{2}}{n-1} \\ &\text { MAPE }=\frac{\sum \frac{\mid \text { Actual }_{t}-\text { Forecast }_{t} \mid}{\text { Actual }_{t}} \times 100}{n} \end{aligned} $$